### v2

@@ Appendix G Analysis for Mirror Prox
Appendix H Omitted Details for Stochastic Convex Optimization
In this section, we present the technical details for Section 5, including further discussions regarding Assumption 4 in Section H.1 and omitted proofs for SMD in Section H.2.
H.1 Comparison of Noise Assumptions
@@
Below, we briefly compare Assumption 4 with commonly used assumptions in stochastic optimization.
@@
When $\sigma(\cdot)$ degenerates into a positive constant, Assumption 4 recovers the uniformly bounded noise condition in the seminal work on $(L_{0},L_{1})$-smoothness (Zhang et al., 2020b), and it is also commonly adopted in other subfields (Koloskova et al., 2023; Zhang et al., 2023; Yu et al., 2024).
@@
If $\sigma(\alpha)=\sigma_{0}+\sigma_{1}\alpha$Assumption 4 recovers the almost sure $(\sigma_{0},\sigma_{1})$-affine noise condition in Liu et al. (2023b); Hong and Lin (2024). Its in-expectation version is also widely adopted in the existing literature (Shi et al., 2021; Jin et al., 2021; Wang et al., 2023a; Jiang et al., 2024).
@@
Another popular setting is that the stochastic gradient has bounded $p$-th moment for some $p\in(1,2]$ (Sadiev et al., 2023; Liu et al., 2024), which includes the uniformly bounded variance condition (Ghadimi and Lan, 2012, 2013; Arjevani et al., 2023) with $p=2$. We underline that the generalized bounded noise is never stronger than this finite moment condition (cf. 27).
@@
This condition is strictly stronger than the finite variance condition and is extremely effective in deriving high-probability bounds (Juditsky et al., 2011; Liu and Zhou, 2024). As demonstrated by 27Assumption 4 can not imply sub-Gaussian noise.
@@
Recently, Liu and Zhou (2025) propose a relaxation of the traditional heavy-tailed noises assumption (e.g., finite moment condition) in the form $\mathbb{E}_{t-1}[\left\|\bm{\epsilon}_{t}\right\|_{2}^{p}]\leq\sigma_{0}^{p}+\sigma_{1}^{p}\left\|\nabla f(\mathbf{x}_{t})\right\|_{2}^{p}$ for some $p\in(1,2],\sigma_{0},\sigma_{1}\in\mathbb{R}_{+}$. The almost sure version of the generalized heavy-tailed noise corresponds to our formulation with $\sigma(\alpha)=\sigma_{0}^{p}+\sigma_{1}^{p}\alpha^{p}$. While Assumption 4 and this condition do not imply each other, both can model heavy-tailed noise environments.
@@
Lastly, we present a simple proof to demonstrate the relative strength of our generalized bounded noise assumption compared to the finite moment condition. A similar argument is presented in Liu and Zhou (2025, Example A.1), who construct a counterexample that satisfies the generalized heavy-tailed noise condition but does not satisfy the finite moment condition.
Proposition 27.
Assumption 4 does not imply the finite moment condition given below:
$\forall t\in\mathbb{N},p\in(1,2],\quad\mathbb{E}_{t-1}[\left\|\bm{\epsilon}_{t}\right\|_{*}^{p}]<+\infty,$ (173)
where the expectation $\mathbb{E}_{t-1}$ is conditioned on the past stochasticity $\left\{\bm{\epsilon}_{s}\right\}_{s=0}^{t-1}$.
Proof H.1.
@@
In this case, the finite moment condition degenerates to the finite variance assumption (Ghadimi and Lan, 2012, 2013; Arjevani et al., 2023). It suffices to show that the variance could be unbounded under Assumption 4 with equality condition in (iii). We proceed with
$\mathbb{E}_{t-1}\left[\left\|\bm{\epsilon}_{t}\right\|_{*}^{2}\right]=$ $\mathbb{E}_{t-1}\left[\sigma_{0}+\sigma_{1}\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\right]^{2}$ (174)
@@
$\mathbb{E}_{t-1}\left[\left\|\bm{\epsilon}_{t}\right\|_{*}^{2}\right]=$ $\sigma_{0}^{2}+4\sigma_{0}\sigma_{1}\mathbb{E}_{t-1}\left[\left\|\mathbf{x}_{t}\right\|_{*}\right]+4\sigma_{1}^{2}\mathbb{E}_{t-1}\left[\left\|\mathbf{x}_{t}\right\|_{*}^{2}\right]$ (175)
@@
which clearly contradicts the finite variance condition. Hence, Assumption 4 is not strictly stronger than the finite moment condition.
H.2 Analysis for Stochastic Mirror Descent
The lemmas in this subsection hold under the conditions in Theorem 12. We do not restate this point for simplicity. First, we present the following two technical lemmas. Lemma 28 is the Hoeffding-Azuma inequality for martingales, which is often used as a standard tool to derive the high probability bound. Lemma 29 is a powerful technique to transform from the convergence behavior from the average-iterate to the last-iterate.
Lemma 28.
@@ H.2 Analysis for Stochastic Mirror Descent
$\forall t\in\mathbb{N},\quad G_{t}:=\left\|\nabla f(\mathbf{x}_{t})\right\|_{*},\ L_{t}:=\ell(2G_{t}),\ \sigma_{t}:=\sigma(G_{t})$ (178)
Lemma 30.
@@
$\eta_{t+1}[f(\mathbf{x}_{t+1})-f(\mathbf{x})]\leq B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x},\mathbf{x}_{t+1})+\eta_{t+1}\langle\bm{\epsilon}_{t},\mathbf{x}-\mathbf{x}_{t}\rangle+\eta_{t+1}^{2}\sigma_{t}^{2}.$ (179)
Proof H.2.
@@
$\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|\leq$ $\eta_{t+1}\left\|\mathbf{g}_{t}\right\|_{*}\leq\eta_{t+1}(\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}+\left\|\bm{\epsilon}_{t}\right\|_{*})$ (181)
@@
where the first step is due to Lemma 17; the last step leverages the non-decreasing property of $\sigma(\cdot)$. Now that (181) holds, we can utilize the local smoothness property in Lemma 3 and we follow the same steps as in (26) and (27) to derive
@@
$+\eta_{t+1}\langle\bm{\epsilon}_{t},\mathbf{x}-\mathbf{x}_{t}\rangle+\eta_{t+1}\langle\bm{\epsilon}_{t},\mathbf{x}_{t}-\mathbf{x}_{t+1}\rangle$
$\leq$ $B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x},\mathbf{x}_{t+1})+\eta_{t+1}\langle\bm{\epsilon}_{t},\mathbf{x}-\mathbf{x}_{t}\rangle-\frac{1}{2}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}$
$+\frac{1}{4}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}+\eta_{t+1}^{2}\left\|\bm{\epsilon}_{t}\right\|_{*}^{2}+\frac{1}{4}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}$
$\leq$ $B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x},\mathbf{x}_{t+1})+\eta_{t+1}\langle\bm{\epsilon}_{t},\mathbf{x}-\mathbf{x}_{t}\rangle+\eta_{t+1}^{2}\sigma_{t}^{2},$
where the second inequality uses Lemma 15 and $\eta_{t+1}\leq 1/(2L_{t})$.
Lemma 31.
@@
$\widetilde{F}_{t}:=30\max\left\{1,\frac{\sigma_{s-1}}{G_{s-1}}\right\}L_{t-1}D^{2}\sqrt{\ln\frac{16}{\delta}}+12D\sigma_{t-1}\sqrt{\ln\frac{16}{\delta}}+4D\sigma_{t-1}\sqrt{\ln\frac{2t^{3}}{\delta}}\ln t,$ (184)
$\widetilde{G}_{t}:=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2\widetilde{F}_{t}\ell(2\alpha)\right\},\quad\forall t\in\mathbb{N}_{+}.$
Under Assumptions 3 and 4, with probability at least $1-\delta$, the following holds simultaneously for all $t\in\mathbb{N}_{+}$:

### v3

@@ Appendix H Analysis for Mirror Prox
Proof of Equation 13.
By Lemma H.2, we have $\left\|\nabla f(\mathbf{y}_{t})\right\|_{*}\leq G<\infty,\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G<\infty$ for all $t\in\mathbb{N}_{+}$. Since the bound for $\nabla f(\mathbf{x}_{0})=\nabla f(\mathbf{y}_{0})$ naturally holds, so we obtain
By Lemma E.1 and (155), we have
As $\eta\leq 1/(2L)$, we set $\beta=1$ in Lemma D.2 to derive
Combine (156) with (158):
Telescoping and taking $\mathbf{x}=\mathbf{x}_{*}$ on both sides of the inequality, we obtain
$\sum_{t=1}^{T}[f(\mathbf{y}_{t})-f^{*}]\leq\frac{B(\mathbf{x}_{*},\mathbf{x}_{0})-B(\mathbf{x}_{*},\mathbf{x}_{T})}{\eta}\leq\frac{B(\mathbf{x}_{*},\mathbf{x}_{0})}{\eta}.$ (161)
Since the output is the average of iterates, i.e., $\overline{\mathbf{y}}_{T}=\sum_{t=1}^{T}\mathbf{y}_{t}/T$, the proof is finished by Jensen’s inequality and the convexity of $f$. ∎
Appendix I Analysis of Stochastic Mirror Descent
The lemmas in this section hold under the conditions of Theorem 4.2. For brevity, we do not restate this point.
Lemma I.1.
Suppose $\eta_{s}\leq\frac{1}{2L}$ and that $\left\|\nabla f(\mathbf{x}_{s})\right\|_{*}\leq G,\left\|\mathbf{g}_{s}\right\|_{*}\leq\frac{G}{\eta_{s+1}L}$ for all $0\leq s\leq t-1$. For any $0<\delta<1$, with probability at least $1-\delta$, for each $t\in[T]$, it holds that
$f(\mathbf{x}_{t})-f(\mathbf{x}_{0})\leq 16\sigma^{2}\log\left(\frac{2}{\delta}\right)\sum_{s=1}^{t}\frac{\eta_{s}^{2}}{\sum_{k=s}^{t}\eta_{k}}.$ (162)
Proof.
This lemma is built upon Liu and Zhou (2025b, Lemma 4.3). Under the conditions of $\left\|\nabla f(\mathbf{x}_{s})\right\|_{*}\leq G,\left\|\mathbf{g}_{s}\right\|_{*}\leq\frac{G}{\eta L}$, in each step $t$ where we utilize $\ell*$-smoothness, it degenerates to the standard $L$-smoothness in Liu and Zhou (2025b). Thus, the proof of Lemma 4.3 in Liu and Zhou (2025b) remains valid. Substituting arbitrary $\mathbf{x}=\mathbf{x}_{0}$, noting that $M=0$ and $1\leq 2\log\left(\frac{2}{\delta}\right)$ finishes the proof. ∎
Lemma I.2 (Bounded suboptimality gaps).
With probability at least $1-\frac{\delta}{2}$, for all $t\in[T]$, we have
$f(\mathbf{x}_{t})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}+\frac{64\sigma}{\sqrt{\delta}}.$ (163)
In the meantime, it holds that
$\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G,\ \left\|\mathbf{g}_{t}\right\|_{*}\leq\frac{G}{\eta_{t+1}L},\ \left\|\boldsymbol{\epsilon}_{t}\right\|_{*}\leq\frac{G}{2\eta_{s+1}L}\quad\forall t-1\in[T].$ (164)
Proof.
The essence underlying this lemma is to enforce the happening of events $\left\{A_{t}\right\}_{t=0}^{T-1}$ and $\left\{B_{t}\right\}_{t=0}^{T-1}$. For the latter, which governs the gradient estimates, we invoke Chebyshev’s inequality under Assumption 4.1:
$\mathbb{P}\left(\bigcup_{t=0}^{T-1}B_{t}\right)\leq\sum_{t=0}^{T-1}\mathbb{P}\left(\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}>\frac{G}{2\eta_{s+1}L}\right)\leq\frac{4\eta^{2}L^{2}\sigma^{2}}{G^{2}}\leq\frac{4\eta^{2}G^{4}\delta}{128^{2}G^{2}}\leq\frac{\delta}{4},$ (165)
where the last inequality is due to $\eta\leq 32/G$ together with the following relation induced by Lemma D.3:
$G^{2}=2LF=2L\left(f(\mathbf{x}_{0})-f^{*}+\frac{64\sigma}{\sqrt{\delta}}\right)\geq\frac{128L\sigma}{\sqrt{\delta}}.$ (166)
On the other hand, for the former $\left\{A_{t}\right\}_{t=0}^{T-1}$, which governs the suboptimality gaps, we seek help from Lemma I.1. We proceed by induction to bound the failure probability of events $\left\{A_{t}\right\}_{t=0}^{T-1}$. For $t=0$, $A_{0}$ trivially holds. By Lemma D.3, we have $\left\|\nabla f(\mathbf{x}_{0})\right\|_{*}\leq G$. Since $B_{0}$ holds with probability at least $1-\frac{\delta}{4T}$ according to (165), we have
$\eta_{1}\left\|\mathbf{g}_{0}\right\|_{*}\leq\eta_{1}\left\|\nabla f(\mathbf{x}_{0})\right\|_{*}+\eta_{1}\left\|\boldsymbol{\epsilon}_{0}\right\|_{*}\leq\frac{G}{2L}+\frac{\eta_{1}G}{2\eta_{1}L}=\frac{G}{L}.$
Then, we invoke Lemma I.1 to derive
$f(\mathbf{x}_{1})-f(\mathbf{x}_{0})\leq\frac{16\eta\sigma^{2}}{\sqrt{T}}\log\left(\frac{8T}{\delta}\right)=\frac{16\sigma}{\sqrt{T}}\sqrt{\log\left(\frac{8T}{\delta}\right)}\quad\text{w.p.}\geq 1-\frac{\delta}{4T},$
[…]
$f(\mathbf{x}_{t})-f(\mathbf{x}_{0})\leq\frac{16\sigma}{\sqrt{T}}\sqrt{\log\left(\frac{8T}{\delta}\right)}\sum_{s=1}^{t}\frac{\eta}{\sqrt{T}(t-s+1)}\leq 16\sigma\left(1+\log(t)\right)\sqrt{\frac{\log\left(\frac{8T}{\delta}\right)}{T}}$
$\leq$ $32\sigma\log(T)\sqrt{\frac{\log\left(\frac{8T}{\delta}\right)}{T}}\leq\frac{64\sigma}{e}\sqrt{\log\left(\frac{8}{\delta}\right)+3}\leq 37\sigma\sqrt{\log\left(\frac{8}{\delta}\right)}\leq\frac{64\sigma}{\sqrt{\delta}},$
[…]
With the above preparations, we are ready to prove the main theorem.
Proof of Theorem 4.2.
By (164) in Lemma I.2, we deduce that throughout the trajectory of (14), we can view the behavior of $\ell*$-smooth functions as the standard smooth functions with effective smoothness parameter $L$. In this way, the conditions for Lemma I.1 are satisfied. According to Eq.(30) in Liu and Zhou (2025b), with probability at least $1-\frac{\delta}{2}$, we have
$f(\mathbf{x}_{T})-f^{*}\leq\frac{4B(\mathbf{x}_{*},\mathbf{x}_{0})}{\sum_{t=1}^{T}\eta_{t}}+4\sigma^{2}\left(1+2\log\left(\frac{4}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}.$
Note that the above relation must build upon the premise of Lemma I.2, i.e., the occurrence of events $\cup_{t=0}^{T-1}A_{t}$ and $\cup_{t=0}^{T-1}B_{t}$. We take a union bound among these events to obtain
$f(\mathbf{x}_{T})-f^{*}\leq\frac{4B(\mathbf{x}_{*},\mathbf{x}_{0})}{\sum_{t=1}^{T}\eta_{t}}+16\sigma^{2}\log\left(\frac{4}{\delta}\right)\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}},\quad\text{w.p.}\geq 1-\delta.$
Noting that $\eta_{t}=\min\left\{\frac{1}{2L},\frac{64}{G},\sqrt{\frac{B(\mathbf{x}_{*},\mathbf{x}_{0})}{\sigma^{2}\log\left(\frac{1}{\delta}\right)\log(T)}}\right\}$, we finally arrive at
$f(\mathbf{x}_{T})-f^{*}\leq$ $O\left(\frac{LB(\mathbf{x}_{*},\mathbf{x}_{0})}{T}+\sigma\sqrt{\frac{B(\mathbf{x}_{*},\mathbf{x}_{0})\log\left(\frac{1}{\delta}\right)\log(T)}{T}}+\frac{GB(\mathbf{x}_{*},\mathbf{x}_{0})}{\sqrt{T}}\right)$
[…]
which holds sufficiently large $T$ and small $\delta$. ∎
@@ Appendix I Analysis of Stochastic Mirror Descent
Appendix J Extension for Stochastic Convex Optimization Under New Noise Model
In this section, we provide new convergence results for SCO under $\ell*$-smoothness and a newly proposed noise model (Assumption J.1). Section J.1 introduces the new noise condition. Section J.2 gives the main result, providing a high-probability anytime convergence guarantee. Section J.3 compares it with other popular noise assumptions, where one can see why this new formulation can be meaningful and more expressive. Section J.4 provides the omitted proof.
J.1 Generalized Bounded Noise Model
We formally introduce the generalized bounded noise assumption as below.
Assumption J.1.
For all $t\in\mathbb{N}$, the stochastic gradients are unbiased:
$\mathbb{E}_{t-1}[\boldsymbol{\epsilon}_{t}]=0,$ (167)
where the expectation $\mathbb{E}_{t-1}$ is conditioned on the past stochasticity $\left\{\boldsymbol{\epsilon}_{s}\right\}_{s=0}^{t-1}$. The noise satisfies the generalized bounded condition w.r.t. a non-decreasing polynomial function $\sigma:\mathbb{R}_{+}\mapsto\mathbb{R}_{++}$:
$\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}\leq\sigma(\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}),\quad\text{a.s.}$ (168)
This generalized bounded noise [^7] The noise is bounded when the gradient is bounded but can become unbounded if the gradient explodes. assumption is inspired by the affine variance condition initially proposed by Bottou et al. (2018), where the noise level depends on the norm of the true gradient and is captured via an affine function. Assumption J.1 encompasses the uniformly bounded noise (Zhang et al., 2020b) and affine noise (Liu et al., 2023b; Hong and Lin, 2024) condition with $\sigma(\alpha)\equiv\sigma$ and $\sigma(\alpha)\equiv\sigma_{0}+\sigma_{1}\alpha$, respectively. Moreover, we underline that Assumption J.1 is never stronger than the commonly assumed finite moment or sub-Gaussian noise condition, with thorough discussions deferred to Section J.3.
For theoretical analysis, we also need to following bounded domain assumption, which is also widely adopted in the context of mirror descent (Zhang et al., 2023, 2024b; Yu et al., 2024; Bai et al., 2025).
Assumption J.2.
The domain $\mathcal{X}$ is bounded in the sense that $\max_{\mathbf{x}\in\mathcal{X}}\psi(\mathbf{x})-\min_{\mathbf{x}\in\mathcal{X}}\psi(\mathbf{x})\leq D^{2}$.
Note that Assumption J.2 does not imply $\sup_{\mathbf{x},\mathbf{y}}B(\mathbf{x},\mathbf{y})\leq D^{2}$, which is used in Eldowa and Paudice (2024, Lemma 9) to simplify the analysis. This stronger assumption is rarely used and can not be satisfied by the simplest instance of Bregman divergence induced by negative entropy functions.
J.2 Theoretical Guarantees
Based on our new noise model, establishing a theoretical guarantee is highly challenging for the sake of the following two quantities that need to be controlled.
Suboptimality gap   Due to the inherent stochasticity, the sequence $\left\{f(\mathbf{x}_{t})-f^{*}\right\}_{t\in\mathbb{N}}$ may become unbounded. Recall from deterministic optimization, we directly link suboptimality gaps to the last-iterate convergence. However, stochastic gradient descent (SGD) (Robbins and Monro, 1951) primarily focuses on the average-iterate, and therefore, such a link does not hold. A fruitful line of work (Shamir and Zhang, 2013; Harvey et al., 2019; Orabona, 2020; Jain et al., 2021) explore the last-iterate convergence of SGD but often remain confined to Euclidean settings, compact domains, or uniformly bounded noise. Our solution is motivated by Orabona (2020), who elegantly converts average-iterate convergence of SGD to last-iterate in-expectation convergence for non-smooth objectives within Euclidean domains. Through a careful analysis of martingales and a refinement of their techniques, we extend the framework to $\ell*$-smooth objectives in non-Euclidean settings. Leveraging this conversion framework, we prove that suboptimality gaps are bounded for all iterations simultaneously with high probability, indicating that the gradients are also bounded for all $t\in\mathbb{N}$. Such “anytime” characteristic is vital for our analysis, since Lemma 2.7 needs to be applied to every round of SMD.
Dual norm of the stochastic gradient   In the context of SCO, we still need to exploit the local smoothness property (Lemma 2.7), which requires the distance between consecutive iterates to be bounded. In the deterministic case, we have $\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|\leq\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}$, and by Lemma 3.4, the boundedness of $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}$ is inferred from finite suboptimality gaps. In contrast, for the stochastic case, it holds that $\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|\leq\left\|\mathbf{g}_{t}\right\|_{*}$, where the challenge lies in bounding $\left\|\mathbf{g}_{t}\right\|_{*}$. Clearly, the self-bounding property in Lemma 3.4 fails to bridge the stochastic gradient and the suboptimality gap. To address this, we leverage the noise model in Assumption J.1 to connect the noise level with the dual norm of the true gradient. As the suboptimality gap is bounded with high probability, the boundedness of the true gradient follows directly, implying that the noise is almost surely bounded via $\sigma(\cdot)$. Finally, we apply a union bound over these events to establish bounds on the stochastic gradient.
Based on the above reasoning, our analysis framework significantly differs from that of Li et al. (2023a, Section 5.2), who utilize a stopping-time technique (Williams, 1991) to show that the two key quantities are bounded before the stopping time $T$ with high probability. However, their method has two limitations: (i) it requires prior knowledge of the total number of iterations $T$; (ii) it lacks a convergence guarantee for the outputs generated before the final round. Moreover, their approach is ill-suited for SMD, as it heavily relies on the Euclidean inner product such that $\langle\mathbf{x},\mathbf{x}\rangle=\left\|\mathbf{x}\right\|_{2}^{2}$. Our main result is presented in Theorem J.3, whose proof is deferred to Section J.4.
Theorem J.3.
Under Assumptions 2.5, J.2, 3.3 and J.1, set
$0<\eta_{t}\leq\min\left\{\frac{D}{\sigma(\left\|\nabla f(\mathbf{x}_{t-1})\right\|_{*})\sqrt{t}},\right.\left.\frac{\min\left\{1,\left\|\nabla f(\mathbf{x}_{t-1})\right\|_{*}/\sigma(\left\|\nabla f(\mathbf{x}_{t-1})\right\|_{*})\right\}}{2\ell(2\left\|\nabla f(\mathbf{x}_{t-1})\right\|_{*})}\right\}.$ (169)
Remark J.4.
Theorem J.3 achieves the classic $\widetilde{O}(1/\sqrt{t})$ anytime convergence rate (Shamir and Zhang, 2013; Zhang et al., 2024a; Bai et al., 2025) in the classic SCO. Note that existing results rely on a deterministic learning rate schedule proportional to $1/\sqrt{t}$. However, this strategy is too coarse to capture the local features of the objective under $\ell*$-smoothness. Consequently, we incorporate local curvature information into the learning rates, akin to the approach in Xie et al. (2024b), who consider online convex optimization under $\ell$-smoothness.
J.3 Comparison of Noise Assumptions
@@
Below, we briefly compare Assumption J.1 with commonly used assumptions in stochastic optimization.
@@
When $\sigma(\cdot)$ degenerates into a positive constant, Assumption J.1 recovers the uniformly bounded noise condition in the seminal work on $(L_{0},L_{1})$-smoothness (Zhang et al., 2020b), and it is also commonly adopted in other subfields (Koloskova et al., 2023; Zhang et al., 2023; Yu et al., 2024).
@@
If $\sigma(\alpha)=\sigma_{0}+\sigma_{1}\alpha$ or $\sigma(\alpha)=\sqrt{\sigma_{0}^{2}+\sigma_{1}^{2}\alpha^{2}}$Assumption J.1 recovers the almost sure $(\sigma_{0},\sigma_{1})$-affine noise condition in Liu et al. (2023b); Hong and Lin (2024) and Attia and Koren (2023), respectively. Its in-expectation version is also widely adopted in the existing literature (Shi et al., 2021; Jin et al., 2021; Wang et al., 2023a; Jiang et al., 2024).
@@
Another popular setting is that the stochastic gradient has bounded $p$-th moment for some $p\in(1,2]$ (Sadiev et al., 2023; Liu et al., 2024), which includes the uniformly bounded variance condition (Ghadimi and Lan, 2012, 2013; Arjevani et al., 2023) with $p=2$. We underline that the generalized bounded noise is never stronger than this finite moment condition (cf. Proposition J.5).
@@
This condition is strictly stronger than the finite variance condition and is extremely effective in deriving high-probability bounds (Juditsky et al., 2011; Liu and Zhou, 2024). As demonstrated by Proposition J.5Assumption J.1 can not imply sub-Gaussian noise.
@@
Recently, Liu and Zhou (2025a) propose a relaxation of the traditional heavy-tailed noises assumption (e.g., finite moment condition) in the form $\mathbb{E}_{t-1}[\left\|\boldsymbol{\epsilon}_{t}\right\|_{2}^{p}]\leq\sigma_{0}^{p}+\sigma_{1}^{p}\left\|\nabla f(\mathbf{x}_{t})\right\|_{2}^{p}$ for some $p\in(1,2],\sigma_{0},\sigma_{1}\in\mathbb{R}_{+}$. The almost sure version of the generalized heavy-tailed noise corresponds to our formulation with $\sigma(\alpha)=\sigma_{0}^{p}+\sigma_{1}^{p}\alpha^{p}$. While Assumption J.1 and this condition do not imply each other, both can model heavy-tailed noise environments.
@@
Lastly, we present a simple proof to demonstrate the relative strength of our generalized bounded noise assumption compared to the finite moment condition. A similar argument is presented in Liu and Zhou (2025a, Example A.1), who construct a counterexample that satisfies the generalized heavy-tailed noise condition but does not satisfy the finite moment condition.
Proposition J.5.
Assumption J.1 does not imply the finite moment condition given below:
$\forall t\in\mathbb{N},p\in(1,2],\quad\mathbb{E}_{t-1}[\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}^{p}]<+\infty,$ (171)
where the expectation $\mathbb{E}_{t-1}$ is conditioned on the past stochasticity $\left\{\boldsymbol{\epsilon}_{s}\right\}_{s=0}^{t-1}$.
Proof.
@@
In this case, the finite moment condition degenerates to the finite variance assumption (Ghadimi and Lan, 2012, 2013; Arjevani et al., 2023). It suffices to show that the variance could be unbounded under Assumption J.1 with equality condition in (iii). We proceed with
$\mathbb{E}_{t-1}\left[\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}^{2}\right]=$ $\mathbb{E}_{t-1}\left[\sigma_{0}+\sigma_{1}\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\right]^{2}$ (172)
@@
$\mathbb{E}_{t-1}\left[\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}^{2}\right]=$ $\sigma_{0}^{2}+4\sigma_{0}\sigma_{1}\mathbb{E}_{t-1}\left[\left\|\mathbf{x}_{t}\right\|_{*}\right]+4\sigma_{1}^{2}\mathbb{E}_{t-1}\left[\left\|\mathbf{x}_{t}\right\|_{*}^{2}\right]$ (173)
@@
which clearly contradicts the finite variance condition. Hence, Assumption J.1 is not strictly stronger than the finite moment condition. ∎
J.4 Proof of Theorem J.3
The lemmas in this subsection hold under the conditions in Theorem J.3. We do not restate this point for simplicity. First, we present the following two technical lemmas. Lemma J.6 is the Hoeffding-Azuma inequality for martingales, which is often used as a standard tool to derive the high probability bound. Lemma J.7 is a powerful technique to transform from the convergence behavior from the average-iterate to the last-iterate.
Lemma J.6.
@@
$\forall t\in\mathbb{N},\quad G_{t}=\left\|\nabla f(\mathbf{x}_{t})\right\|_{*},\ L_{t}=\ell(2G_{t}),\ \sigma_{t}=\sigma(G_{t})$ (176)
Lemma J.8.
@@
$\eta_{t+1}[f(\mathbf{x}_{t+1})-f(\mathbf{x})]\leq B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x},\mathbf{x}_{t+1})+\eta_{t+1}\langle\boldsymbol{\epsilon}_{t},\mathbf{x}-\mathbf{x}_{t}\rangle+\eta_{t+1}^{2}\sigma_{t}^{2}.$ (177)
Proof.
@@
$\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|\leq$ $\eta_{t+1}\left\|\mathbf{g}_{t}\right\|_{*}\leq\eta_{t+1}(\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}+\left\|\boldsymbol{\epsilon}_{t}\right\|_{*})$ (179)
@@
where the first step is due to Lemma E.1; the last step leverages the non-decreasing property of $\sigma(\cdot)$. Now that (179) holds, we can utilize the local smoothness property in Lemma 2.7 and we follow the same steps as in (64) and (65) to derive
@@
$+\eta_{t+1}\langle\boldsymbol{\epsilon}_{t},\mathbf{x}-\mathbf{x}_{t}\rangle+\eta_{t+1}\langle\boldsymbol{\epsilon}_{t},\mathbf{x}_{t}-\mathbf{x}_{t+1}\rangle$
$\leq$ $B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x},\mathbf{x}_{t+1})+\eta_{t+1}\langle\boldsymbol{\epsilon}_{t},\mathbf{x}-\mathbf{x}_{t}\rangle-\frac{1}{2}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}$
$+\frac{1}{4}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}+\eta_{t+1}^{2}\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}^{2}+\frac{1}{4}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}$
$\leq$ $B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x},\mathbf{x}_{t+1})+\eta_{t+1}\langle\boldsymbol{\epsilon}_{t},\mathbf{x}-\mathbf{x}_{t}\rangle+\eta_{t+1}^{2}\sigma_{t}^{2},$
where the second inequality uses Lemma D.2 and $\eta_{t+1}\leq 1/(2L_{t})$. ∎
Lemma J.9.
@@
[…]
$\widetilde{G}_{t}=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2\widetilde{F}_{t}\ell(2\alpha)\right\},\quad\forall t\in\mathbb{N}_{+}.$
Under Assumptions 3.3 and J.1, with probability at least $1-\delta$, the following holds simultaneously for all $t\in\mathbb{N}_{+}$: